OpenStudy (anonymous):

The equation of line AB is y = 2x + 4. Write an equation of a line parallel to line AB in slope-intercept form that contains point (3, -2).

OpenStudy (anonymous):

Select one: a. y = -2x + 4 b. y= -1/2x -1/2 c. y= 1/2x - 7/2 d. y= 2x-8

OpenStudy (anonymous):

@Nurali

OpenStudy (anonymous):

There is only one line parallel to AB. It is y=2x-8 and this line goes thru (3,-2)

OpenStudy (anonymous):

so the answer is D

OpenStudy (anonymous):

yes